# Algorithmic resolution of crowd-sourced moderation on X in polarized settings across countries

### v1

@@ Abstract
Social platforms increasingly transition from expert fact-checking to crowd-sourced moderation, with X pioneering this shift through its Community Notes system, enabling users to collaboratively moderate misleading content. To resolve conflicting moderation, Community Notes learns a latent ideological dimension and selects notes garnering cross-partisan support. As this system, designed for and evaluated in the United States, is now deployed worldwide, we evaluate its operation across diverse polarization contexts. We analyze 1.9 million moderation notes with 135 million ratings from 1.2 million users, cross-referencing ideological scaling data across 13 countries. Our results show X’s Community Notes effectively captures each country’s main polarizing dimension but fails by design to moderate the most polarizing content, posing potential risks to civic discourse and electoral processes.
Keywords: social media, misinformation, moderation, crowd-sourced fact-checking.
Social media platforms have become ubiquitous sources for the consumption of news content  (?), where users may also participate in content creation and dissemination. A persistent challenge in these platforms is the dissemination of misinformation  (?, ?, ?), prompting platforms to navigate a trade-off between freedom of expression on the one hand, and content moderation on the other  (?, ?). While expert fact-checking has been shown to successfully decrease the extent to which people agree with misleading claims  (?, ?, ?), it faces significant limitations in practice: i) professional fact-checking operations cannot scale to meet the volume of content requiring verification  (?, ?), ii) their efforts may be disconnected from online attention  (?), iii) moderation may be perceived as having partisan bias by the public  (?), and iv) they may suffer from low visibility  (?). Additionally, professional moderation may backfire by further “entrenching” polarized individuals  (?). As an alternative, crowd-sourced moderation of content (i.e., systems in which regular platform users signal misinformed or potentially misleading content) has demonstrated that panels of regular users can effectively identify misinformation  (?, ?, ?, ?).
Motivated by these findings, X (formerly Twitter) piloted a crowd-sourced fact-checking system in the United States in early 2021  (?), later deployed worldwide in 2023 under the name Community Notes  (?). This system enables self-selected contributors to attach contextual notes to posts they consider “misinformed or potentially misleading”. Other contributors can then rate these notes based on their subjective assessment of the content and its context. A core challenge for the platform is moderating polarized notes, for which consensus on usefulness may not emerge between raters across the ideological spectrum.
@@
Computationally, X first converts ratings to numerical values: 1 for Helpful, 0.5 for Somewhat Helpful, and 0 for Not Helpful  (?). Then, to predict these empirical ratings, the system infers several parameters: note latent ideologies ($\theta_{n}$), rater latent ideologies ($\theta_{r}$), individual rater leniency ($\beta_{r}$), the tendency of notes to be rated helpful regardless of raters’ ideologies ($\beta_{n}$), and a global helpfulness baseline ($\beta_{0}$). These parameters are estimated, from the ratings alone, to best predict the rating $\eta$ that rater $r$ will assign to note $n$ through the predictive formula:
$\hat{\eta}_{rn}=\beta_{0}+\beta_{n}+\beta_{r}+\theta_{n}\cdot\theta_{r}.$
@@
Consequently, a note rated Helpful by diverse raters (i.e., with $\theta_{r}$ values spread across positive and negative values) will have a higher $\beta_{n}$ value and lower absolute value $|\theta_{n}|$ when compared to notes rated Helpful by only one side of the spectrum. X’s Community Notes service uses these parameters to algorithmically resolve conflicting ratings by selecting notes with the highest $\beta_{n}$ values  (?) and displays them under the post, visible to all users on the platform  (?).
@@
Results from X’s pilot phase in the United States revealed that Community Notes selected for display garnered broad ideological appeal  (?) and that users exposed to such notes were less likely to share the posts and less inclined to agree with the posts’ misleading claims. Independent research found high agreement between Community Notes and expert fact-checker classifications  (?). Additionally, Allen et al.  (?) showed that political partisanship was highly predictive of contributors’ ratings in the US, with raters preferentially challenging content from those with whom they disagree politically. Following X’s Community Notes global deployment in 2023, subsequent studies corroborated these findings, documenting reduced sharing of flagged content  (?) and accurate sourcing of credible references for notes addressing COVID-19 vaccination misinformation  (?). Moreover, moderation resulting from the Community Notes service was shown to be less prone to backfire effects that entrench individuals  (?, ?) and generated greater user trust than warnings issued by expert fact-checkers  (?).
X’s Community Notes system was developed and tested in the United States  (?, ?), where political disagreement can be effectively modeled using a single ideology dimension  (?, ?). However, the relevance of the internal representation of note and rater ideology —central to the algorithmic moderation process in Community Notes— remains largely unexplored in more general political settings. In countries where political competition is organized along multiple ideological dimensions  (?, ?, ?), what form of polarization is captured by the latent ideology in X’s Community Notes, and how this relates to the performance of this crowd-sourced moderation systems, remains an open question. In this study, we examine X’s Community Notes usage and outcomes across 13 countries and compare the latent ideological space inferred by X’s algorithm with established political dimensions derived from follower networks  (?) and calibrated with political survey data  (?).
@@
To compare the ideological scale inferred by X’s system with established ideological dimensions across countries, we rely on the method developed by Ramaciotti et al.  (?), which adapts Barberá’s  (?) methodology to multidimensional settings. Specifically, we identify X accounts following Members of Parliament (MPs) from 23 countries across five continents, retaining those that followed at least 3 MPs and were followed by at least 25 accounts, following Barberá’s protocol  (?, ?, ?). We then infer, through Correspondence Analysis, a multidimensional latent space arranging MPs and their followers by homophily (see details in SM.SE.1).
For the analysis of Community Notes latent ideology parameter, we consider the position of X accounts along $\delta_{1}$, the first principal component of the Correspondence Analysis, which best explains which MPs they follow. Additionally, we map the positions of accounts onto interpretable political dimensions of the Global Party Survey  (?) —an expert survey positioning parties along several dimensions endowed with a reference frame making them comparable across countries— namely i) the Left-Right dimension, a leading dimension of polarization relevant for studies of misinformation in social media  (?, ?, ?), and ii) the Anti-Elite dimension, which measures criticism toward elites and institutions and has been proposed as an alternative explanation for the dissemination of misinformation online  (?, ?, ?). When comparing latent ideologies $\theta_{n}$ and $\theta_{r}$ with Left-Right and Anti-Elite dimensions, we focus on 13 countries where at least a thousand X accounts with inferred political positions have authored posts with proposed Community Notes. These countries represent a geographically diverse sample (see SM.SE.3): United States, United Kingdom, Japan, Spain, France, Brazil, Canada, Germany, Argentina, Israel, Australia, Poland, and Mexico.
@@ Results
Finally, we observe in Figures 2.A & 2.B that, across analyzed countries, Community Notes are requested and proposed under posts authored by accounts throughout the political spectrum. Of the non-deleted posts that received proposed Community Notes and have authors with inferred ideological leanings, 58% are from right-leaning accounts in the United States, 54% in Japan, and 53% in France (see left-right leaning distribution per country in SM.SE.3).
@@
We now examine the status assigned by X’s Community Notes algorithm to proposed notes.
Of all proposed Community Notes arguing the original post is “misinformed or potentially misleading”, only 11.97% garner sufficient favorable ratings from “diverse” contributors to attain Helpful Status and be made visible on X. The remaining notes are not publicly displayed, either because consensus determined them to be Not Helpful (2.81%; hereinafter, achieving Not Helpful Status), or because they lacked sufficient ratings from diverse contributors to reach a conclusive status (85.22%). 17.1% of posts for which at least 5 users requested Community Notes ultimately received a proposed note that reached Helpful Status. Between January and March 2025, half of the notes that achieved Helpful Status did so within 4h47min from note submission and 15h17min after the original post publication; see distribution in Figure 1.B.
@@
Similarly, we observe that notes discussing politically divisive issues struggle to garner consensus and achieve Helpful Status across each analyzed country, while those reporting scams and other misleading practices commonly do (see SM.SD.2 for country-specific lists of over- and under-represented terms in notes with Helpful Status). For instance, in the United States, only 6.0% of notes discussing presidential candidates and political parties, 5.5% of those addressing the Israel-Palestine conflict, and 11.7% of those covering the Russia-Ukraine war reach Helpful Status. In contrast, 39.0% of notes reporting scams, fraud, and impersonation achieve this status (see SM.SD.2 for a curation of keywords used).
@@
In the United States, the single-dimensional latent ideological space learned by X’s Community Notes system closely aligns with users’ Left-Right political leanings, notes with $\theta_{n}>0$ appear under posts published by Left-leaning accounts, while notes with $\theta_{n}<0$ are associated with Right-leaning accounts (Fig. 3.A). In particular, 95.0% of notes associated with posts authored by Republican members of Congress have $\theta_{n}<0$, while 94.6% of those associated with Democratic members of Congress have $\theta_{n}>0$ (see the average latent ideologies $\theta_{n}$ of notes associated with key political figures across multiple countries in SM.SG.1). Complementarily, contributors with $\theta_{r}<0$ tend to rate notes as Helpful when they appear under Left-leaning accounts but rate notes as Not Helpful when they appear under Right-leaning accounts (Fig. 3.B), and the opposite pattern holds for raters with $\theta_{r}>0$ (Fig. 3.C), aligning with the observations made by Allen et al. during the pilot program  (?). In the United States, the Left-Right political leaning of users predicts the sign of the latent ideologies $\theta_{n}$ of the notes associated with their posts, with an AUC of $0.822$. Additionally, news outlets cited in Community Notes align ideologically with contributor leanings; we found a Pearson correlation of $\rho=0.744$ ($p<10^{-4}$) between Ad Fontes media bias ratings  (?) and the latent ideology $\theta_{r}$ of contributors citing these media outlets as sources in their notes (see SM.SG.1.1).
@@
Figure 3: (A to C) Distribution of Left-Right leaning of X accounts (A) segmented by the sign of the community note latent ideology $\theta_{n}$, (B) rated helpful/not-helpful by raters with negative latent ideology $\theta_{r}<0$, (C) rated helpful/not-helpful by raters with positive latent ideology $\theta_{r}>0$. (D to G) X accounts positioned in the (Left-Right; Anti-Elite) 2D plane, segmented by the majority sign of the latent ideology $\theta_{n}$ of their associated Community Notes (red for $\theta_{n}<0$ and blue for $\theta_{n}>0$). The direction that best separates notes by majority outcome, learned through logistic regression, is displayed as $CN$ along with the corresponding AUC score for sign prediction based on account positions in the 2D plane (standard deviation from 10-fold cross-validation). The direction structuring each country’s political landscape $\delta_{1}$ is projected onto the plane.
The latent ideology computed by the Community Notes model ($\theta_{n}$ and $\theta_{r}$) shows varying degrees of alignment with the Left-Right and the Anti-Elite dimensions computed with ideology scaling and political surveys. Figures 3.D to 3.G map X users with proposed Community Notes along the Left-Right and Anti-Elite dimensions in the United Kingdom, Japan, France, and Brazil. The latent dimension learned by X’s system ($CN$ in Fig. 3) —along which notes and raters are positioned— aligns closely with each country’s most structuring direction ($\delta_{1}$ in Fig. 3) in this Left-Right/Anti-Elite plane. We quantify this alignment by measuring how well the position of a user along $\delta_{1}$ predicts the sign of the latent ideology $\theta_{n}$ of notes associated with their posts. AUC values range from 0.850 in Poland to 0.729 in Israel, averaging $0.808\pm 0.037$ across 13 countries. Complementarily, user positions in the Left-Right/Anti-Elite plane predict the sign of the latent ideology $\theta_{n}$ of notes with an AUC of $0.813\pm 0.035$ across all 13 countries (see breakdown for all countries in our study in SM.SG.2).
We find that the direction learned by X’s system to structure content veracity disputes ($CN$ in Fig. 3) aligns with the Left-Right dimension but cannot always be reduced to it, as anticipated by recent literature  (?, ?). Instead, it may present additional alignments with attitudes towards elites and institutions (and, to a lesser degree, possibly additional issue dimensions not considered in our study), as highlighted by the cases of Japan and Brazil (Fig. 3.F & 3.G).

### v2

# Community Notes undermoderate polarizing content by design creating risks in electoral processes

@@ Abstract
Community Notes (CNs) of X enables users to collaboratively moderate misleading content. To resolve conflicting moderation, CNs infers a latent ideological dimension and selects notes garnering cross-partisan support. As this system is now deployed worldwide, we evaluate its operation across diverse polarization contexts. We analyze all 1.9 million moderation notes receiving 135 million ratings by March 2025, cross-referencing ideological scaling data on 13 countries. Our results show that the CNs algorithm effectively captures the main polarizing dimensions across countries, surfacing notes that garner cross-partisan support. This also means that, by design, CNs systematically under-moderate polarizing content. We analyze notes relating to four recent elections in the US (2024), the UK (2024), France (2024) and Germany (2025) and demonstrate that they are systematically under-moderated when compared to other notes, posing potential risks to civic discourse and electoral processes.
Social media platforms have become ubiquitous sources for the consumption of news content (?), where users may also participate in content creation and dissemination. A persistent challenge in these platforms is the dissemination of misinformation (?, ?, ?), prompting platforms to deploy means for content moderation (?, ?). While expert fact-checking has been shown to successfully decrease the extent to which people agree with misleading claims (?, ?, ?), this approach faces significant limitations in practice: i) professional fact-checking operations cannot scale to meet the volume of content requiring verification (?, ?), ii) their efforts may be disconnected from online attention (?), iii) moderation may be perceived as having partisan bias by the public (?), and iv) they may suffer from low visibility (?). Additionally, professional moderation may backfire by further “entrenching” polarized individuals (?). On the other hand, research has shown that crowd-sourcing fact-checking assessments from regular users can effectively identify misinformation, with politically balanced laypeople ratings correlating strongly with those of professional fact-checkers (?, ?, ?, ?).
Motivated by these findings, X (formerly Twitter) piloted a crowd-sourced fact-checking system in the United States in early 2021 (?), later deployed worldwide in 2023 under the name Community Notes (?). This system enables self-selected contributors to attach contextual notes to posts they consider “misinformed or potentially misleading”. Other contributors can then rate the helpfulness of these notes based on their subjective assessment of the content and its context. A core challenge for the platform is moderating polarizing content, for which consensus on usefulness may not emerge between raters across the ideological spectrum (e.g., when raters preferentially seek to moderate posts by political adversaries).
@@
Computationally, X first converts ratings given to notes to numerical values: 1 for Helpful, 0.5 for Somewhat Helpful, and 0 for Not Helpful (?). Then, to predict these empirical ratings, the system infers several parameters: note latent ideologies ($\theta_{n}$), rater latent ideologies ($\theta_{r}$), individual rater leniency ($\beta_{r}$), the tendency of notes to be rated helpful regardless of raters’ ideologies ($\beta_{n}$), and a global helpfulness baseline ($\beta_{0}$). These parameters are estimated, from the ratings alone, to best predict the rating $\eta$ that rater $r$ will assign to note $n$ through the predictive formula $\hat{\eta}_{rn}=\beta_{0}+\beta_{n}+\beta_{r}+\theta_{n}\cdot\theta_{r}$.
@@
Consequently, a note rated Helpful by diverse raters (i.e., with $\theta_{r}$ values spread across positive and negative values) will have a higher $\beta_{n}$ value and lower absolute value $|\theta_{n}|$ when compared to notes rated Helpful by only one side of the spectrum. X’s Community Notes service uses these parameters to algorithmically resolve conflicting ratings by selecting notes with the highest $\beta_{n}$ values (?) and displaying them under the post, visible to all users on the platform (?).
@@
Results from X’s pilot phase in the United States revealed that notes selected for display garnered broad appeal across ideological positions (?) and that users exposed to such notes were less likely to share the posts and less inclined to agree with the posts’ misleading claims. Independent research found high agreement between these algorithmically selected notes and expert fact-checker classifications (?). Additionally, Allen et al. (?) showed that political partisanship was highly predictive of contributors’ ratings in the US, with raters preferentially challenging content from those with whom they disagree politically. Following X’s Community Notes global deployment in 2023, subsequent studies corroborated these findings, documenting reduced sharing of flagged content (?) and accurate sourcing of credible references for notes addressing COVID-19 vaccination misinformation (?). Moreover, moderation resulting from the Community Notes service was shown to be less prone to backfire effects that entrench individuals (?, ?) and generated greater user trust than warnings issued by expert fact-checkers (?).
X’s Community Notes system was developed and tested in the United States (?, ?), where online polarization is traditionally modeled using a single ideology dimension (?, ?). However, the relevance of this single-dimensional representation of note and rater ideology —central to the algorithmic moderation process in Community Notes— remains largely unexplored in other national settings that might be structured along several ideology dimensions (?, ?, ?). Whether the Community Notes model captures relevant political dimensions in different countries has a direct impact in how the service functions, and it determines what type of polarizing content will not be algorithmically moderated due to lack of sufficient agreement among raters. In this study, we examine X’s Community Notes usage and algorithmic outcomes across all countries in which the service is available. Leveraging ideology scaling of follower networks (?) calibrated with political survey data (?), we compare the latent ideological parameter inferred by Community Note’s algorithm with the main dimensions of political competition in 13 countries. Our results show that in all 13 countries we study, the Community Notes model reliably identifies the main ideological divide that structures political communication networks on X. This remains true even after accounting for a second dimension related to misinformation: trust in elites and institutions (?, ?). This implies that notes are proposed, rated, and surfaced as Helpful across the ideological spectrum. However, because this system only surfaces notes that receive broad cross-ideological agreement, content that strongly activates polarization is systematically under-moderated. We demonstrate the importance of this intrinsic limitation by showing that the share of notes related to elections is consistently lower than for other topics during four major elections in our study period: the 2024 U.S. Presidential election, the 2024 U.K. general election, the 2024 French legislative elections, and the 2025 German federal elections. While Community Notes offers several important benefits, this limitation poses a significant risk for information integrity during electoral processes.
@@
To examine the ideological scale inferred by X’s system, we rely on the method developed by Ramaciotti et al. (?), which adapts Barberá’s (?) methodology to multidimensional settings. Specifically, we identify X accounts following Members of Parliament (MPs) from 23 countries across five continents, retaining those that followed at least 3 MPs and were followed by at least 25 accounts, following Barberá’s protocol (?, ?, ?). We then infer, through Correspondence Analysis, a multidimensional latent space arranging MPs and their followers by homophily (see details in SM.SE.1).
For the analysis of Community Notes latent ideology parameter, we consider the position of X accounts along $\delta_{1}$, the first principal component of the Correspondence Analysis, which best explains which MPs they follow. Additionally, we map the positions of accounts onto interpretable political dimensions of the Global Party Survey (?) —an expert survey positioning parties along several dimensions endowed with a reference frame making them comparable across countries— namely i) the Left-Right dimension, a leading dimension of polarization relevant for studies of misinformation in social media (?, ?, ?), and ii) the Anti-Elite dimension, which measures criticism toward elites and institutions. In contrast to results regarding online misinformation in the US (?), the degree of negative attitudes towards and institutions has been shown to play a role in the probability of disseminating misinformation (?, ?). When comparing latent ideologies $\theta_{n}$ and $\theta_{r}$ with Left-Right and Anti-Elite dimensions, we focus on 13 countries where at least a thousand X accounts with inferred political positions have authored posts with proposed Community Notes. These countries represent a geographically diverse sample (see SM.SE.3): United States, United Kingdom, Japan, Spain, France, Brazil, Canada, Germany, Argentina, Israel, Australia, Poland, and Mexico.
@@ Results
Finally, we observe in Figures 3.A & 3.B that, across analyzed countries, Community Notes are requested and proposed under posts authored by accounts throughout the political spectrum. Of the non-deleted posts that received proposed Community Notes and have authors with inferred ideological leanings, 58% are from right-leaning accounts in the United States, 54% in Japan, and 53% in France (see left-right leaning distribution per country in SM.SE.3).
@@
We now examine the status assigned by X’s Community Notes algorithm to proposed notes. Of all proposed Community Notes arguing the original post is “misinformed or potentially misleading”, only 11.97% garner sufficient favorable ratings from “diverse” contributors to attain Helpful Status and be made visible on X. The remaining notes are not publicly displayed, either because consensus determined them to be Not Helpful (2.81%; hereinafter, achieving Not Helpful Status), or because they lacked sufficient ratings from diverse contributors to reach a conclusive status (85.22%). 17.1% of posts for which at least 5 users requested Community Notes ultimately received a proposed note that reached Helpful Status. Between January and March 2025, half of the notes that achieved Helpful Status did so within 4h47min from note submission and 15h17min after the original post publication; see distribution in Figure 1.B.
@@
Figure 2: Fraction of notes reaching Helpful Status per country during elections. The analysis compares election-related notes with the overall corpus (in the same period, ranging from one month before to one week after the election) during the 2024 United States presidential, 2024 United Kingdom general, 2024 France legislative, and 2025 Germany federal elections. Error bars represent 95% confidence intervals, estimated by bootstrapping over the keywords used to identify election-related notes.
Similarly, we observe that notes discussing politically divisive issues struggle to garner consensus and achieve Helpful Status across each analyzed country, while those reporting scams and other misleading practices commonly do (see SM.SD.2 for country-specific lists of over- and under-represented terms in notes with Helpful Status). To demonstrate this, we analyzed notes related to four elections taking place during the period of observation of our study. Figure 2 shows a comparison between the fraction of notes reaching Helpful Status for election-related notes and for overall notes in the same period, for the 2024 United States presidential (6.9% vs 14.7%), 2024 United Kingdom general (8.2% vs 10.9%), 2024 France legislative (12.8% vs 16.8%), and 2025 Germany federal elections (8.6% vs 11.9%). For each selected election, consensus is reached less frequently on election-related community notes compared to overall notes published during the same period. For comparison, 39% of notes reporting scams, frauds, and impersonations reached Helpful Status. See SM.SD.2 for a curation of keywords used to identify notes relating to elections and other topics.
@@
In the United States, the single-dimensional latent ideological space learned by X’s Community Notes system closely aligns with users’ Left-Right political leanings, notes with $\theta_{n}>0$ appear under posts published by Left-leaning accounts, while notes with $\theta_{n}<0$ are associated with Right-leaning accounts (Fig. 4.A). 95.0% of notes associated with posts authored by Republican members of Congress have $\theta_{n}<0$, while 94.6% of those associated with Democratic members of Congress have $\theta_{n}>0$ (see the average latent ideologies $\theta_{n}$ of notes associated with key political figures across multiple countries in SM.SG.1). Contributors with $\theta_{r}<0$ tend to rate notes as Helpful when they appear under Left-leaning accounts but rate notes as Not Helpful when they appear under Right-leaning accounts (Fig. 4.B), and the opposite pattern holds for raters with $\theta_{r}>0$ (Fig. 4.C), aligning with the observations made by Allen et al. during the pilot program (?). In the United States, the Left-Right political leaning of users predicts the sign of the latent ideologies $\theta_{n}$ of the notes associated with their posts, with an AUC of $0.822$. Additionally, news outlets cited in Community Notes align ideologically with contributor leanings; we found a Pearson correlation of $\rho=0.744$ ($p<10^{-4}$) between Ad Fontes media bias ratings (?) and the latent ideology $\theta_{r}$ of contributors citing these media outlets as sources in their notes (see SM.SG.1.1).
Figure 4: (A to C) Distribution of Left-Right leaning of X accounts (A) segmented by the sign of the community note latent ideology $\theta_{n}$, (B) rated helpful/not-helpful by raters with negative latent ideology $\theta_{r}<0$, (C) rated helpful/not-helpful by raters with positive latent ideology $\theta_{r}>0$. (D to G) X accounts positioned in the (Left-Right; Anti-Elite) 2D plane, segmented by the majority sign of the latent ideology $\theta_{n}$ of their associated Community Notes (red for $\theta_{n}<0$ and blue for $\theta_{n}>0$). The direction that best separates notes by majority outcome, learned through logistic regression, is displayed as CNCN along with the corresponding AUC score for sign prediction based on account positions in the 2D plane (standard deviation from 10-fold cross-validation). The direction structuring each country’s political landscape $\delta_{1}$ is projected onto the plane.
@@
The latent ideology computed by the Community Notes model ($\theta_{n}$ and $\theta_{r}$) shows varying degrees of alignment with the Left-Right and the Anti-Elite dimensions computed with ideology scaling and political surveys. Figures 4.D to 4.G map X users with proposed Community Notes along the Left-Right and Anti-Elite dimensions in the United Kingdom, Japan, France, and Brazil. The latent dimension learned by X’s system ($CN$ in Fig. 4) —along which notes and raters are positioned— aligns closely with each country’s most structuring direction ($\delta_{1}$ in Fig. 4) in this Left-Right/Anti-Elite plane. We quantify this alignment by measuring how well the position of a user along $\delta_{1}$ predicts the sign of the latent ideology $\theta_{n}$ of notes associated with their posts. AUC values range from 0.850 in Poland to 0.729 in Israel, averaging $0.808\pm 0.037$ across 13 countries. Complementarily, user positions in the Left-Right/Anti-Elite plane predict the sign of the latent ideology $\theta_{n}$ of notes with an AUC of $0.813\pm 0.035$ across all 13 countries (see breakdown for all countries in our study in SM.SG.2).
We find that the direction learned by X’s system to structure content veracity disputes (direction $CN$ in Fig. 4) aligns with the Left-Right dimension but cannot always be reduced to it, as anticipated by recent literature (?, ?). Instead, it may present additional alignments with attitudes towards elites and institutions (and, to a lesser degree, possibly additional issue dimensions not considered in our study), as highlighted by the cases of Japan and Brazil (Fig. 4.F & 4.G).